OpenStudy (anonymous):

2i + j = a(i + j) +b(i - j) solve for a and b

OpenStudy (anonymous):

First use distributive property on right hand side z*(x + y) = z*x + z*y So what do you get a *(i + j) = --------------? & b*(i - j) = ------------------?

OpenStudy (anonymous):

Are you there ?